OpenStudy (anonymous):

Hi I know how to get a percentage of a number but how can I figure out what a number is from a percentage. For instance 337 is 106%. How would I find out what the original number is?

OpenStudy (anonymous):

Let the original number be X then 106/100 * X = 337 X= (337*100)/106

Parth (parthkohli):

Wait... what? @Harkirat is back!

OpenStudy (anonymous):

translate directly in to math 337 is the number 337 "is" means "=" 106% means 1.06 "of" means \(\times\) some number pick a variable, i will use \(x\) we get \[337=1.06x\] solve by \[x=337\div 1.06\]

Parth (parthkohli):

\[337 = 1.06x \]

OpenStudy (anonymous):

Yeah logged in after many days☺

Parth (parthkohli):

Satellite wins at winning.

OpenStudy (anonymous):

Thanks Harkirat! I got the answer, when rounded 317.

OpenStudy (anonymous):

U r welcome....